OpenStudy (anonymous):

Use the intermediate value theorem to determine whether the following the equation has a solution or not. If so, then use a graphing calculator to solve the equation. cosx=x Please help me understand this!

OpenStudy (anonymous):

From \(-\pi\) to \(\pi\), \(\cos(x)\) ranges from \(-1\) to \(1\) to \(-1\) again.

OpenStudy (anonymous):

They probably want you to use this fact.

OpenStudy (anonymous):

To be more specific, \(\cos(0) = 1\) and \(\cos(\pi)=0\).

OpenStudy (nincompoop):

this is so vague that you have the freedom to choose what your value for c is going to be

OpenStudy (anonymous):

Then, what would x be approx.?

OpenStudy (anonymous):

Indeed, it is very vague and I think it's vague-ness is throwing me off.

OpenStudy (anonymous):

Okay, here is another idea. Let \(f(x)=\cos(x)-x\). Use intermediate value theorem to find out if there is any interval where \(f(x)=0\).

OpenStudy (nincompoop):

that makes it easier, @wio

OpenStudy (nincompoop):

but you should provide the interval as well

OpenStudy (anonymous):

Same as before. \(f(0) = 1\) and \(f(\pi) = -\pi\). We know \(f(x)\) is continuous so it mus go through \(0\).

OpenStudy (nincompoop):

thank you, that's all folks

OpenStudy (anonymous):

Did you use a calculator to come to that conclusion or was that knowledge you knew before hand?

OpenStudy (anonymous):

I actually understand trigonometry, so I know certain values of \(\cos(x)\).

OpenStudy (nincompoop):

wio is a human calculator

OpenStudy (anonymous):

I understand trig functions but this problem seems like it requires more than one answer and I'm still so confused.

zepdrix (zepdrix):

So you've used the IVT to determine that `yes` a solution exists. The next part is just to use a calculator of some sort to approximate the solution. If you don't have a graphing calculator, use Wolfram as a nice resource.